8. A university student painter is considering the purchase of a new air compressor and paint gun to replace an old paint sprayer. Suppose the CCA rates are 20% and ent, respectively. The two new items cost $8,500 and have a useful life of four years, at which time they can be sold for $1,100. The old paint sprayer can be sold now for $275 and be scraped for $150 in four years. The student is very confident that operating revenue will increase 25% for the old equipment and new equipm 2 annually by $7,600. If tax rate is 24% and the required rate of return is 19%, should the purchase be made? What is the maximum price of the new equipment that makes the student to replace the old equipment?
Answer #1

Let's calculate this in four different parts and then pool together the result to get the final output.

Part -1

Initial outlay at t = 0

AN = Cost of new equipment = 8,500

AO = Sale Price of old equipment = 275

Hence, Net Initial Investment = NII = net cash outflow at t=0 will be = AN - AO = 8,500 - 275 = 8,225

Part -2

Salvage Value at the end of year 4,

SN = Salvage Value of new equipment = 1,100

SO = Salvage value of old equipment = 150

Hence, Net salvage value, NSV = net cash inflow at t=4 will be = SN - SO = 1,100 - 150 = 950

PV(NSV) = PV of Net Salvage Value = NSV / (1 + r)N where r = discount rate = 19% and N = life of the project = 4 years

Hence, PV(NSV) = 950 / (1 + 19%)4 = 473.74

Part - 3

Post tax incremental operating profit = (1 - T) x Incremental operating profit

Incremental operating profit = incremental revenue in this case = 7,600

Tax rate = t = 24%

Hence, post tax incremental operating profit = (1-24%) x 7,600 = 5,776 per year for next four years

This is an annuity for N = 4 years as this incremental profit occurs every year for next four years

Hence, PV (Incremental post tax income) = Incremental post tax income x Sum of PV factors @ 19% over 4 years

(1-(1+r) (1-(1 + 0.19)-4) = = 5.776 × 5,776 × 15, 240.47 0.19

Part - 4

PV(NCCATS) = Present value of Net CCA tax shield

PV (CCATS) = present value of CCA tax shield for an asset is given by the formula:

PV (CCTS) (r d)(1 r) r+d(1r)N

where T = tax rate, d = CCA depreciation rate, A = original asset cost, S = Salvage value, N = life, r= discount rate

For the new equipment,

PV (CCTS) (r d)(1 r) r+d(1r)N

0.24 × 0.25 × 8.500 × (1 +0.5 × 0.19) 0.190.25) (1 0.19) 0.24 × 0.25 × 1, 100 0.190.2510.19)4 991.76

For the old equipment,

PV (CCTS) (r d)(1 r) r+d(1r)N

0.24 × 0.20 × 275 × (1 +0.5 × 0.19) 0.19 +0.20 (10.19) 0.24 × 0.20 × 150 0.19 0.20) (10.19)4 21.94

PV (NCCTS) = PV(CCTS) of new equipment - PV (CCTS) of old equipment = 991.76 - 21.94 = 969.82

NPV of the replacement decision = - NII + PV(NSV) + PV(Post tax incremental profit) + PV(NCCTS) = - 8,225 + 473.74 + 15,240.47 + 969.82 = $ 8,459.03

Since it's positive NPV, the purchase of new equipment should be undertaken.

Maximum price that one should pay = Cost of new equipment that makes NPV equal to zero. Let's say the incremental cost is C. Then S should be such that

C - PV (CCATS) on C = NPV

Hence, we need to solve the following equation:

C- (r d)(1 r)

0.24 × 0.25 × C × (1 +0.5 × 0.19) 0.190.25) (1 0.19) C- = 8.459.03

Hence C = 9,672.74

Hence the maximum price = Original price + incremental price = AN + C = 8,500 + 9,672.74 =  $18,172.74